# IDs for AI Systems

### v1

@@ 8 Related Work
Besides real-world examples of IDs and work on documenting AI system that we highlighted in Section 1, we discuss two additional lines of related work.
A growing line of work is concerned with building digital infrastructure to structure how AI systems, especially agents (Chan et al. 2023; Shavit et al. 2023; Gabriel et al. 2024), behave and interact. Such infrastructure can have a variety of goals. Patil et al. (2024) build a runtime for LLM agents to enable human validation and reversal of actions. Marro (2024) sketches a protocol for communication between LLMs; the protocol relies on documents to specify structured rules, with natural language as a fallback. (Sun et al. 2023) argue that decentralized commitment devices will be necessary to allow agents to coordinate with each other. IDs are a form of infrastructure that enable informed decision-making about AI interactions.
The technical implementations we assess take inspiration from several Internet and security protocols. HTTPS (Fielding, Nottingham, and Reschke 2022) uses digital certificates (Cooper et al. 2008) to allow users to verify the identity of an accessed website. If the certificate is valid, a user’s browser uses the information contained in the certificate to establish a secure connection (Rescorla 2018). Websites must obtain such certificates from CAs (Cooper et al. 2008), who verify the identity of the website owner and issue the certificate upon successful identification. Identity verification depends upon public key cryptography (Barnes et al. 2019), which allows a party to prove ownership of a given, public identifier, otherwise known as a public key. Let’s Encrypt is a non-profit CA which automates the process of issuing certificates (Encrypt 2024). HTTPS everywhere was a browser plug-in that forced the usage of the HTTPS version of websites whenever it was available; this functionality is now default in modern web browsers (Hancock 2021). To design further digital infrastructure for AI systems, other internet and security protocols may be a fruitful source of inspiration.
@@ 9 Conclusion
To inform crucial decisions about AI interactions, we proposed a framework for creating IDs for AI systems. IDs can vary in their attributes, how accessible they are to various parties, and to what extent they are verifiable. There could be significant demand for IDs from governments, service providers, and users, particularly when systems engage in high-stakes tasks. These actors also have means to incentivize ID usage. ID implementation seems feasible for deployers, but implementation in decentralized settings will require further study.
More research is required to understand and address the potential risks of IDs. First, security and privacy risks should be addressed, such as the potential for certificate authorities to become compromised. Second, it is unclear under what circumstances IDs should be accompanied by user identification. Third, the broader societal consequences of introducing IDs require further study.
Limited experimentation with IDs seems justified given their potential benefits and risks. IDs seem most warranted in high-stakes settings, where certain actors could experiment with incentives for ID use. For instance, service providers that provide plugins in potentially high-stakes settings (e.g., plug-ins that allow AI systems to make financial transactions) could experiment with rate-limiting instances without IDs, while still allowing such instances to access services. Deployers could begin to implement the technical infrastructure for IDs.
IDs have facilitated essential functions across a variety of domains. Incident investigation, allocation of liability, and establishment of trust would be more difficult without IDs for products, organizations, and software systems. As AI systems become increasingly prevalent, deficiencies in such functions could result in more severe consequences for society. If implemented well, IDs for AI systems could enable mechanisms to avert such consequences and navigate this emerging world.
@@ Appendix A A More Detailed Definition of an Instance
Appendix B Maintaining the Accessibility of IDs
Suppose a primary party $P$ receives an output and ID from an AI system. If it is desirable for secondary parties to see the ID, $P$ may have additional responsibilities. Some potential situations follow.
@@ Appendix B Maintaining the Accessibility of IDs
Users: Users could receive outputs with attached IDs, whether as a watermark or metadata. To counter inadvertent removal of IDs, deployers could add labels to inform users. For example, Facebook and Instagram automatically add labels indicating AI origin when a user shares a photo generated with the Meta AI feature (Bickert 2024). To counter intentional removal of IDs, better watermarking techniques may be required. Yet, it remains unclear how effective watermarking may ultimately be (Zhang et al. 2023).

### v3

@@ 8 Related Work
A growing line of work is concerned with building digital infrastructure to structure how AI systems, especially agents (Chan et al., 2023; Shavit et al., 2023; Gabriel et al., 2024), behave and interact. Patil et al. (2024) build a runtime for LLM agents to enable human validation and reversal of actions. Marro (2024) sketches a protocol for communication between LLMs; the protocol relies on documents to specify structured rules, with natural language as a fallback. (Sun et al., 2023) argue that decentralized commitment devices will be necessary to allow agents to coordinate with each other. Our work builds upon the preliminary discussion of IDs in Chan et al. (2024); Shavit et al. (2023).
Our treatment of verifiability takes inspiration from several Internet and security protocols. HTTPS (Fielding et al., 2022) uses digital certificates (Cooper et al., 2008) to allow users to verify the identity of an accessed website. If the certificate is valid, a user’s browser uses the information contained in the certificate to establish a secure connection (Rescorla, 2018). Websites must obtain such certificates from CAs (Cooper et al., 2008), who verify the identity of the website owner and issue the certificate upon successful identification. Identity verification depends upon public key cryptography (Barnes et al., 2019), which allows a party to prove ownership of a given, public identifier, otherwise known as a public key. Let’s Encrypt is a non-profit CA which automates the process of issuing certificates (Encrypt, 2024). HTTPS everywhere was a browser plugin that forced the usage of the HTTPS version of websites whenever it was available; this functionality is now default in modern web browsers (Hancock, 2021). To design further digital infrastructure for AI systems, other internet and security protocols could be a fruitful source of inspiration (Sporny et al., 2022; 2024; Microsoft, 2023).
There is much work on verifying the provenance of digital entities. Securing software supply chains (Ziv et al., 2024; Springett et al., 2020; Souppaya et al., 2022) has become an increasing priority for organizations and governments (House, 2023). A related subject is data provenance, which focuses on tracking the origin and modification of pieces of data in varied contexts, such as production warehouses, scientific research, and environmental protection (Buneman et al., 2001; Simmhan et al., 2005). Labeling and verifying the provenance of AI-generated outputs has recently received much attention (C2PA, 2023; Zhang et al., 2023; Gleichauf & Geer, 2024).
@@ 9 Conclusion
To inform crucial decisions about AI interactions, we proposed a framework for creating IDs for AI systems. IDs can vary in their attributes, how accessible they are to various parties, and to what extent they are verifiable. IDs could be useful in several scenarios, including for shutting down malfunctioning systems, verifying certifications, and incident investigation. There could be significant demand for IDs from governments, service providers, and users, particularly when systems engage in high-stakes tasks. These actors also have means to incentivize ID usage. ID implementation seems feasible for deployers, but implementation in decentralized settings will require further study.
More research is required to understand and address the potential risks of IDs. First, more work is required to understand what information is appropriate to attach to IDs, and in what contexts. Second, we need to understand better how IDs should be used. Third, when ID verifiability fails, there should be ways to limit the impacts of such failure. Fourth, the broader societal consequences of introducing IDs require further study.
Limited experimentation with IDs seems justified given the balance of potential benefits and risks. IDs seem most warranted in settings where AI systems could have a large impact upon the world, such as in making financial transactions or contacting real humans. in these cases, service providers could experiment with incentives for ID use. For instance, they could rate limit instances without IDs, while still allowing such instances to access services. Deployers could experiment with ID implementation.
IDs have facilitated essential functions across a variety of domains. Incident investigation, allocation of liability, and establishment of trust would be more difficult without IDs for products, organizations, and software systems. As AI systems become increasingly prevalent, deficiencies in such functions could make it more difficult to manage risks from AI systems. If implemented well, IDs for AI systems could enable mechanisms to navigate this emerging world.
@@ Appendix A A More Detailed Definition of an Instance
Appendix B Maintaining Access to IDs
Suppose a primary party $P$ receives an output and ID from an AI system. If it is desirable for secondary parties to access the ID, $P$ may have additional responsibilities. Some potential situations follow.
@@ Appendix B Maintaining Access to IDs
Users: Users could receive outputs with attached IDs, whether as a watermark or metadata. To counter inadvertent removal of IDs, deployers could add labels to inform users. For example, Facebook and Instagram automatically add labels indicating AI origin when a user shares a photo generated with the Meta AI feature (Bickert, 2024). To counter intentional removal of IDs, better watermarking techniques may be required. Yet, it remains unclear how effective watermarking may ultimately be (Zhang et al., 2023; Gleichauf & Geer, 2024).